OpenStudy (sw33thearts):

The sum of the squares of 2 consecutive negative integers is 41. What are the numbers? Which of the following equations is the result of using the factoring method to solve the problem? (n - 5)(n + 4) = 0 (n + 5)(n - 4) = 0 (n + 5)(n + 4) = 0

OpenStudy (sw33thearts):

dont just give the answer

OpenStudy (faiqraees):

Lets take our first integer to be x and second integer (x-1) Thus the squares will be x² and (x-1)² respectively. The sum of the squares will be x² + (x-1)² Which equals to 41 and now we can set an equation x² + (x-1)² = 41 Now simplify this equation

OpenStudy (ineedhelplz):

Let the two numbers be x and x+1 Their squares are x^2 and x^2 + 2x + 1 The sum of their squares is 2x^2 + 2x + 1 = 41 Subtract 41: 2x^2 + 2x - 40 = 0 Divide by 2: x^2 + x - 20 = 0 Factor: (x + 5) (x - 4) = 0 < Here is the answer. Use the zero product principle twice: x = -5, 4 x+1 = -4, 5 So the two consecutive negative numbers are -5 and -4.